OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 4x + 3y = 2 3x + 2y = 1

OpenStudy (anonymous):

multiply equation (1) by 2 and equation (2) by 3 thus 8x+6y=2 9x+6y=3 (2) subtract (1) x=1 finaly y= -2/3